What is the intensity of sound at 122 dB?

Compare it to that of a whisper at 20 dB?

(both answers in W/m^2 units)
Answer #1

intensity level = 10*log(I/Io)

Io = threshold intensity = 10^-12 w/m^2

for intensity level = 122 dB


122 = 10*log(I1/10^-12)

I = 1.58 w/m^2

+++++++++


for intensity level = 20 dB


20 = 10*log(I2/10^-12)

I2 = 10^-10 w/m^2